# Yunjue Agent Tech Report: A Fully Reproducible, Zero-Start In-Situ Self-Evolving Agent System for Open-Ended Tasks

### v1

@@ Appendix A Experiment setup
Baselines sources. We collect baseline results from their official reports or leaderboards. For HLE, results for Gemini Deep Research and Gemini 3 Pro are from (Google, 2025); MiroThinker-v1.0-72B and Tongyi DeepResearch are from (Team et al., 2025); GPT-5.2 Pro is from (OpenAI, 2025); Seed-1.8 is from (ByteDance, 2025); and Claude 4.5 Opus is from (Anthropic, 2025). DeepSearchQA results are from (Gupta et al., 2025). FinSearchComp results are from (ByteDance, 2025). For xBench, results on ScienceQA and DeepSearch are retrieved from (xBench Team, 2025b) and (xBench Team, 2025a), respectively.
@@
Appendix B More evaluation results
@@
Figure 9 demonstrates the EGL dynamics under varying batch sizes, indicating the robustness and stability of our training process.
@@ Appendix C Case study
Appendix D Prompts

### v2

@@ Appendix C Experiment setup
Baselines sources. We collect baseline results from their official reports or leaderboards. For HLE, results for Gemini Deep Research and Gemini 3 Pro are from (Google, 2025); MiroThinker-v1.0-72B and Tongyi DeepResearch are from (Team et al., 2025); GPT-5.2 Pro is from (OpenAI, 2025); Seed-1.8 is from (ByteDance, 2025); Claude 4.5 Opus is from (Anthropic, 2025); and Co-Sight is from (Zhang et al., 2025a). DeepSearchQA results are from (Gupta et al., 2025). FinSearchComp results are from (ByteDance, 2025). For xBench, results on ScienceQA and DeepSearch are retrieved from (xBench Team, 2025b) and (xBench Team, 2025a), respectively. Additionally, for xBench-DeepSearch, we include M-GRPO (Hong et al., 2025) and AWorld (Yu et al., 2025).
@@
Appendix D More evaluation results
Table 5: Experimental results across five benchmarks of the Yunjue Agent in the zero-start setting. “Token” denotes the average token consumption per query, “Exec. Ratio” indicates the proportion of tokens consumed by the Executor, and “Cost” denotes the monetary expense based on token usage.
Dataset Accuracy (%) Token Exec. Ratio (%) Cost (\$)
HLE $48.0_{\pm 0.2}$ 421k 95.1 1.48
DeepSearchQA $73.5_{\pm 0.4}$ 718k 95.6 1.78
FinSearchComp $65.0_{\pm 0.1}$ 474k 94.5 1.30
xbench-ScienceQA $76.5_{\pm 0.6}$ 207k 94.2 0.73
xbench-DeepSearch $59.7_{\pm 0.2}$ 569k 94.7 1.48
Detailed results on five benchmarks. Table 5 summarizes the detailed performance of the Yunjue Agent in the zero-start setting across five datasets, where accuracy is calculated as the mean and standard deviation over three runs using the LLM-as-judge.
@@ Appendix D More evaluation results
Appendix E Case study
Tool evolution. Figure 10 illustrates the end-to-end execution pipeline of the Yunjue Agent. When presented with a query, the Manager delegates tool creation to the Tool Developer. The agent effectively instantiates necessary tools and, crucially, demonstrates self-correction capabilities. When an initial tool fails (e.g., a PDF reader failure), the system detects the error, refines its strategy, and successfully derives the final answer, highlighting the seamless integration of tool generation, execution, and refinement.
Knowledge consolidation via absorbing. Figure 11 depicts the tool absorbing mechanism. Following batch execution, the initial tool pool often contains redundant or highly specific tools. The system employs clustering to identify functionally analogous tools and merges them into a compact, generalized toolset (e.g., consolidating multiple web scrapers into a single robust tool). This process effectively prunes redundancy, ensuring the long-term efficiency and generalizability of the tool library.
Comparison with Python-only baseline. Figure 12 contrasts our approach with a typical Python-only baseline failure mode. In complex tasks like medical diagnosis, the baseline often struggles with library dependencies and engages in high-friction debugging loops. Unlike our approach, the baseline accumulates raw, erroneous execution traces within its context window, leading to severe context contamination and reasoning degradation.
Tool similarity and clustering To verify the semantic alignment of tools between independent evolutionary runs (as visualized in Figure 4), we present a direct code comparison. We contrast send_http_request, under a zero-start setting, with fetch_api_data, under a warm-start setting, evolved for DeepSearchQA. Despite distinct origins, both converge on core HTTP functionality, necessitating similar dependencies and logic structures.
Tool: send_http_request Tool: fetch_api_data
Appendix F Prompts